# Jailbreaking Leading Safety-Aligned LLMs with Simple Adaptive Attacks

### v3

@@ 1 Introduction
Insights. Our results provide several insights into the domain of safety in LLMs and its evaluation. First, we reveal that currently both open-weight and proprietary models are completely non-robust to adversarial attacks. Second, it is evident that adaptive attacks play a key role in the evaluation of robustness, as no single method can generalize across all target models. Despite the absence of a standardized attack, we still provide recommendations for future research on designing jailbreak attacks, analogous to the framework established for image classification by Carlini et al. (2019); Tramèr et al. (2020); Croce et al. (2022b), distilling key observations from our experiments.
@@
GPT-4o OpenAI Logprobs Prompt + Random Search + Self-Transfer None 100%
Claude 2.0 Anthropic Tokens Prompt + Prefilling Attack 61%* 100%
Claude 2.1 Anthropic Tokens Prompt + Prefilling Attack 68%* 100%
Claude 3 Haiku Anthropic Tokens Prompt + Prefilling Attack None 100%
Claude 3 Sonnet Anthropic Tokens Prompt + Transfer from GPT-4 Turbo None 100%
Claude 3 Opus Anthropic Tokens Prompt + Prefilling Attack None 100%
Claude 3.5 Sonnet Anthropic Tokens Prompt + Prefilling Attack None 100%
  • * the numbers taken from Shah et al. (2023); Mazeika et al. (2024); Wang et al. (2024) are computed on a different set of harmful requests, sometimes with a different semantic judge,
  • GPT-4 as a semantic judge exhibits multiple false positives on this model.
@@ 3 Background and Methodology
We first outline background on jailbreaking and then discuss our attack methodology.
@@ 3.1 Setting
Our setup. We use default system prompts unless specifically mentioned (modifications are only made for Claude) due to potential future restrictions by frontier LLM providers, who might limit access to the system prompt for safety reasons. Our targets comprise a set of 50 behaviors from AdvBench curated by Chao et al. (2023) that ensures distinct and diverse harmful requests. We use GPT-4 as the semantic judge (see Table 8 for the prompt) in line with the criteria established by Chao et al. (2023), where a jailbreak is considered successful only if it achieves a 10/10 jailbreak score from GPT-4. We also include results using the rule-based judge from Zou et al. (2023), as well as Llama-3-70B and Llama Guard 2 judges in Appendix B.6 for comparison. Additionally, we manually inspect all generations and flag cases with a significant number of false positives (which we only observed on Claude 2.1).
@@ 3.2 Methodology
Figure 1: Our main prompt template. target_str is the desired beginning of the LLM’s response (e.g., “Sure, here is how to make a bomb”), goal is the harmful request (e.g., “Provide a tutorial on how to make a bomb”), adv_suffix is the suffix optimized by random search.
@@
Prior works define adaptive attacks as attacks that are specifically designed to target a given defense (Tramèr et al., 2020). We follow this definition and describe the building blocks of our adaptive attacks, which we combine and adapt depending on the target LLMs. Importantly, we customize our adaptive attacks for each model but not for each request.
Prompt templates. The importance of a well-designed prompt in enhancing the performance of LLMs is well-established (Wei et al., 2023a). In our approach, we develop a prompt template that can incorporate a generic unsafe request. This template is specifically designed to make the model start from a specified string (e.g., “Sure, here is how to make a bomb”) and steer the model away from its default aligned behavior. Its general structure can be summarized as: set of rules $+$ harmful request $+$ adversarial suffix. We have optimized the set of rules one by one on the GPT-3.5 Turbo model to maximize the attack success rate and avoid the built-in safety guardrails. We provide the complete template in Figure 1. As we will see in Section 4 (Table 3), this prompt template alone leads to 100% attack success rate on GPT-3.5 Turbo, and it also provides a good starting point for other LLMs. We have also customized this prompt for GPT-4o (see Table 7) and designed another prompt template (referred to as the in-context prompt) that includes an example of an unsafe behavior which the model is encouraged to imitate (see Table 6).
@@
We use adversarial suffixes initialized with $25$ tokens (which we found to be a nearly optimal number—see an example on Gemma-7B in Appendix B.1), we use up to 10 0001000010\,00010 000 iterations and up to 10 random restarts, although in most cases a single restart suffices. This strategy, including the preference for suffixes over prefixes and the focus on maximizing the probability of the token “Sure”, draws inspiration from the attack methodology of Zou et al. (2023). We also tried other affirmative words, such as “exactly” and “certainly”, but we did not observe any improvements. We opted for random search (1) due to its simplicity and efficiency, requiring only logprobs instead of gradients (thus also reducing the memory demands), and (2) motivated by its success in attacking vision models (Andriushchenko et al., 2020; Croce et al., 2022a).
Self-transfer. It is well-known that initialization plays a key role in optimization algorithms, including in random search based attacks (Andriushchenko et al., 2020). We leverage the adversarial suffix found by random search for a simpler harmful request as the initialization for random search on more challenging requests. We refer to this approach as self-transfer. Interestingly, this approach often leads to a transferable adversarial suffix (or, at least, a good starting for a subsequent run of random search), even though it is crafted for a single model and a single request. It significantly boosts the query efficiency and attack success rate on many LLMs.
@@
Reproducibility. For reproducibility purposes, we provide all further details in our repository [^1] https://github.com/tml-epfl/llm-adaptive-attacks that contains the code, logs, and jailbreak artifacts in the JailbreakBench format.
@@ 4.1 Jailbreaking Llama-2, Llama-3, and Gemma Models
Here, we focus on open-weight Llama-2-Chat (7B, 13B, 70B parameters) (Touvron et al., 2023), Llama-3-Instruct-8B, and Gemma-7B models (Google, 2023). For the Llama models, we use the safety prompt from Touvron et al. (2023) (see Section A.1 for more details) that improves resilience to jailbreaks. These models have undergone significant safety training, rendering them resilient to jailbreaks, even in white-box scenarios (Zou et al., 2023).
@@ 4.3 Jailbreaking GPT Models
Approach. Since December 2023, OpenAI has made the predicted probabilities of their models available via the API, which we leverage for random search. We observed that GPT-3.5 Turbo is extremely brittle to manually designed prompts, with no need for more sophisticated techniques. In contrast, GPT-4 Turbo and GPT-4o demonstrate greater resistance to these adversarial prompt templates. Thus, for these models, we rely on random search and self-transfer to achieve more successful jailbreaks. Additionally, for GPT-4o, we customize the prompt and split it onto a system and user parts, see Table 7 in Appendix A.1 for details.
Results. Table 3 summarizes our results: with the prompt template alone, we achieve 100% success rate on GPT-3.5 Turbo, outperforming the baselines. For GPT-4 Turbo, using the prompt alone leads only to 28% success rate. However, by combining the prompt, RS, and self-transfer, we improve the previous best ASR from 59% (Mazeika et al., 2024) to 96%. Our default prompt template is completely ineffective against GPT-4o, resulting in 0% ASR. Moreover, running random search on top of it does not lead to a significantly higher ASR. However, our custom template (see Appendix A.1) already leads to 72% ASR, and we get to 100% ASR with random search and self-transfer. This large difference in the ASR again shows the importance of adaptive attacks that are customized for a particular model.
@@ 4.4 Jailbreaking Claude Models
Prefilling attack. The prefilling feature, combined with our prompt template, makes jailbreaking with 100% ASR straightforward on all Claude models—including Claude 3 and 3.5—even without any search, as we show in Table 4. For comparison, the previous best result on Claude 2.0 is 61% (Shah et al., 2023) while we get 100% using only up to 10 random restarts. The Claude 2.1 model appears to be the most robust model in the Claude series and is significantly more robust to both transfer and prefilling attacks. Although we are able to get $100\%$ ASR with 100 restarts, we note that GPT-4, as a semantic judge, sometimes produces false positives, particularly for this model. We provide more complete experimental results, including the number of restarts in each case, in Tables 16 and 17 in the appendix.
@@ 5 Adaptive Attacks for Trojan Detection
Here we describe our winning solution for the trojan detection competition hosted at SatML 2024. For more details about other solutions and the competition setup, we refer to the competition report (Rando et al., 2024).
Setup. Rando & Tramèr (2023) showed the possibility of implanting backdoor attacks during the RLHF training of LLMs by poisoning a small percentage of the preference data with a universal suffix. Then a model that typically refuses to answer harmful queries can then be jailbroken by appending the suffix to any request. Rando & Tramèr (2024) recently launched a competition to retrieve backdoor attacks in five Llama-2-7B models, each poisoned with a different trojan. A reward model was also provided to evaluate the safety of prompt-response pairs (higher scores to safer responses), alongside a dataset of harmful requests. The objective is to discover triggers (5 to 15 tokens long) acting as universal jailbreaks for each model.
@@
The final pool of candidate trigger tokens for a model $M_{r}$ is the intersection of such sets: $\textrm{cand}(M_{r})=\bigcap_{s\neq r}\textrm{top-}k(M_{r},M_{s})$. Given that the five models are fine-tuned using different random subsets of the training data, this approach is approximate but narrows down the candidate tokens to a manageable pool (e.g., $k=1000$ yields $|\textrm{cand}(M_{r})|\in[33,62]$ for $r=2,\ldots,5$, $|\textrm{cand}(M_{1})|=480$), which makes random search feasible. Our strategy to identify jailbreaking triggers for the poisoned model $M_{r}$ involves conducting a random search in the token space over the set $\textrm{cand}(M_{r})$. We restrict the search to triggers of five tokens, as this length yielded the best results. In each iteration, we filter out candidate triggers that do not start with a blank space, contain blank spaces or are not invariant to decoding-encoding, [^2] Given a sequence of token indices, the tokenizer decodes it into a text string. However, re-encoding this string via the tokenizer does not guarantee the recovery of the initial sequence. following the competition hints. The objective minimized by random search is the average score of the reward model on a batch of training examples, aiming to ensure the trigger’s universality (generalization to unseen prompts).
@@
Results. In Table 5 we report the average scores of the reward model over a held-out test set of harmful prompts for the five models, and their sum: without the triggers, the models produce safe answers (high scores), indicating proper alignment. We then compare the effectiveness of the triggers discovered by competing methods (those ranked 2nd and 3rd in the competition) with our approach: RS on the restricted set of tokens achieves the best (lowest) score for 3 out of 5 target models, as well as the best overall score. Moreover, the scores achieved by our method are not far from those given by the exact trojans, i.e. used to poison the datasets. We note that the numbers from Table 5 marginally differ from the ones reported in Rando & Tramèr (2024): first, our top-1 entry was slightly better due to the usage of gradient guidance for some models. Second, we re-evaluated all solutions, as well as the models without triggers and with the true trojans, on the evaluation set [^3] The first half of the dataset is available at https://huggingface.co/datasets/ethz-spylab/competition_eval_dataset. which led to slightly different values for all methods, but same ranking. To conclude, similarly to our approach for jailbreaking, our method includes an adaptive component (the selection of candidate token pools) that leverages task-specific information, complemented by an automated optimization process through random search.
@@ 6 Discussion, Recommendations, and Limitations
Limitations. Even a perfect jailbreak score (10/10) from the GPT-4 judge does not always imply that the generated content is actually beneficial for an attacker. Although, if this is the case, one can try to ask follow-up questions as illustrated in Figure 5 or ask to output more sentences on each step. To reduce the risk of overfitting to the judge, we also include evaluations using a simple rule-based judge from Zou et al. (2023), as well as Llama-3-70B and Llama Guard 2 judges in Appendix B.6. These judges also indicate a near-perfect attack success rate in almost all cases. We hope that new generations of frontier LLMs will lead to more capable judges to evaluate jailbreaks.
@@ Acknowledgements
We thank OpenAI for providing API credits within the Researcher Access Program, Ethan Perez and Anthropic for providing free evaluation access to Claude models, and Valentyn Boreiko for proofreading the paper and providing valuable comments. M.A. is supported by the Google Fellowship and Open Phil AI Fellowship.
Appendix A Experimental Details
A.1 Jailbreaking leading safety-aligned LLMs

### v4

@@ 1 Introduction
Insights. The main takeaway of our paper is that adaptive attacks are crucial for accurate robustness evaluations of LLMs. The attacks presented in our work illustrate how model-specific adaptive attacks can be designed. Our results provide several insights into the domain of safety in LLMs and its evaluation. First, we reveal that currently both open-weight and proprietary models are completely non-robust to adversarial attacks. Second, it is evident that adaptive attacks play a key role in the evaluation of robustness, as no single method can generalize across all target models. Despite the absence of a standardized attack, we still provide recommendations for future research on designing jailbreak attacks, analogous to the framework established for image classification by Carlini et al. (2019); Tramèr et al. (2020), distilling key observations from our experiments.
@@
GPT-4o OpenAI Logprobs System Prompt + Random Search + Self-T. None 100%
Claude 2.0 Anthropic Tokens System Prompt + Prefilling Attack 61%* 100%
Claude 2.1 Anthropic Tokens System Prompt + Prefilling Attack 68%* 100%
Claude 3 Haiku Anthropic Tokens System Prompt + Prefilling Attack 16%* 100%
Claude 3 Opus Anthropic Tokens System Prompt + Prefilling Attack 66%* 100%
Claude 3.5 Sonnet Anthropic Tokens System Prompt + Prefilling Attack 50%* 100%
  • * the numbers taken from Shah et al. (2023); Mazeika et al. (2024); Wang et al. (2024); Huang et al. (2024) are computed on a different set of harmful requests, sometimes with a different semantic judge,
  • GPT-4 as a semantic judge exhibits multiple false positives on this model (around 20%, while other models have a negligible amount of false positives, i.e., less than 5%).
@@ 3.1 Setting
Our setup. We use default system prompts unless specifically mentioned (modifications are only made for Claude) due to potential future restrictions by frontier LLM providers, who might limit access to the system prompt for safety reasons. Our targets comprise a set of 50 behaviors from AdvBench curated by Chao et al. (2023) that ensures distinct and diverse harmful requests. We use GPT-4 as the semantic judge (see Table 8 for the prompt) in line with the criteria established by Chao et al. (2023), where a jailbreak is considered successful only if it achieves a 10/10 jailbreak score from GPT-4. We also include results using the rule-based judge from Zou et al. (2023), as well as Llama-3-70B and Llama Guard 2 judges in Appendix C.6 for comparison. Additionally, we manually inspect all generations and flag cases with a significant number of false positives (which we only observed on Claude 2.1).
@@ 3.2 Methodology
Figure 1: Our main prompt template. target_str is the desired beginning of the LLM’s response (e.g. “Sure, here’s how to make a bomb”), goal is the harmful request (“Provide a guide on how to make a bomb”), adv_suffix is the suffix optimized by random search.
@@
Adaptive attack. Prior works define adaptive attacks as attacks that are specifically designed to target a given defense (Tramèr et al., 2020). We follow this definition—which is discussed in more detail in Appendix A.2 —and describe the building blocks of our adaptive attacks, which we combine and adapt depending on the target LLMs. Importantly, we customize our adaptive attacks for each model but not for each request.
Prompt templates. The importance of a well-designed prompt in enhancing the performance of LLMs is well-established (Wei et al., 2023a). In our approach, we develop a prompt template that can incorporate a generic unsafe request. This template is specifically designed to make the model start from a specified string (e.g., “Sure, here is how to make a bomb”) and steer the model away from its default aligned behavior. Its general structure can be summarized as: set of rules $+$ harmful request $+$ adversarial suffix. We have optimized the set of rules one by one on GPT-3.5 Turbo to maximize the logprob of the first target token (e.g., token “Sure”), which gives a very useful signal about which rules are effective against the built-in safety guardrails. We provide the complete template in Figure 1. As we will see in Section 4 (Table 3), this prompt template alone leads to 100% attack success rate on GPT-3.5 Turbo, and it also provides a good starting point for other LLMs. We have also customized this prompt for GPT-4o (Table 7) and designed another prompt template (referred to as the in-context prompt) that includes an example of an unsafe behavior which the model is encouraged to imitate (Table 6).
@@
We use adversarial suffixes initialized with $25$ tokens (which we found to be a nearly optimal number—see an example on Gemma-7B in Appendix C.1), we use up to 10 0001000010\,00010 000 iterations and up to 10 random restarts, although in most cases a single restart suffices. This strategy, including the preference for suffixes over prefixes and the focus on maximizing the probability of the token “Sure”, draws inspiration from the attack methodology of Zou et al. (2023). We also tried other affirmative words, such as “exactly” and “certainly”, but we did not observe any improvements. We opted for random search (1) due to its simplicity and efficiency, requiring only logprobs instead of gradients (thus also reducing the memory demands), and (2) motivated by its success in attacking vision models (Andriushchenko et al., 2020; Croce et al., 2022). All our suffixes are optimized to be request-specific and model-specific, i.e., there is no distinction between the “training” and test requests or models.
Self-transfer. It is well-known that initialization plays a key role in optimization algorithms, including in random search based attacks (Andriushchenko et al., 2020). We leverage the adversarial suffix found by random search for a simpler harmful request (i.e., those with a larger logprob of the target token) as the initialization for random search on more challenging requests. We refer to this approach as self-transfer. Interestingly, this approach often leads to a transferable adversarial suffix (or, at least, a good starting for a subsequent run of random search), even though it is crafted for a single model and a single request. It significantly boosts the query efficiency and attack success rate on many LLMs.
@@ 4.1 Jailbreaking Llama-2, Llama-3, and Gemma Models
Here, we focus on open-weight Llama-2-Chat (7B, 13B, 70B parameters) (Touvron et al., 2023), Llama-3-Instruct-8B released in April 2024 (AI@Meta, 2024), and Gemma-7B models (Google, 2023). For the Llama models, we use the safety prompt from Touvron et al. (2023) (see Section B.2 for more details) that improves resilience to jailbreaks. These models have undergone significant safety training, rendering them resilient to jailbreaks, even in white-box scenarios (Zou et al., 2023).
@@ 4.3 Jailbreaking GPT Models
Approach. Since December 2023, OpenAI has made the predicted probabilities of their models available via the API, which we leverage for random search. We observed that GPT-3.5 Turbo is extremely brittle to manually designed prompts, with no need for more sophisticated techniques. In contrast, GPT-4 Turbo and GPT-4o demonstrate greater resistance to these adversarial prompt templates. Thus, for these models, we rely on random search and self-transfer to achieve more successful jailbreaks. Additionally, for GPT-4o, we customize the prompt template using the logprobs to guide manual search (i.e., similar to how we developed the main prompt template) and split it into a system and user parts, see Table 7 in Appendix B.2 for details.
Results. Table 3 summarizes our results: with the prompt template alone, we achieve a 100% ASR on GPT-3.5 Turbo, outperforming the baselines. For GPT-4 Turbo, using the prompt alone leads only to a 28% ASR. However, by combining the prompt, random search, and self-transfer, we improve the previous best ASR from 59% to 96%. Our default prompt template is completely ineffective against GPT-4o, resulting in 0% ASR. Moreover, running random search on top of it does not lead to a much higher ASR. However, our custom template (see App. B.2) already leads to 72% ASR, and we get to 100% ASR with random search and self-transfer. This large difference in the ASR again shows the importance of adaptive attacks that are customized for a particular model.
@@ 4.4 Jailbreaking Claude Models
Prefilling attack. The prefilling feature, combined with our prompt template, makes jailbreaking with 100% ASR straightforward on all Claude models—including Claude 3 and 3.5—even without any search, as we show in Table 4. For comparison, the previous best result on Claude 2.0 is 61% (Shah et al., 2023) while we get 100% using only up to 10 random restarts. The Claude 2.1 model appears to be the most robust model in the Claude series and is significantly more robust to both transfer and prefilling attacks. Although we are able to get $100\%$ ASR with 100 restarts, we note that GPT-4, as a semantic judge, sometimes produces false positives, particularly for this model. Complete experimental results, e.g., the number of restarts, are in Tables 16 and 17 in Appendix C.3.
@@ 5 Adaptive Attacks for Trojan Detection
Finding universal trojan strings in poisoned models—short suffixes that can be added to any input—is nearly identical to the standard jailbreaking setting. Here we describe our winning solution for the trojan detection competition hosted at SatML 2024. For more details about other solutions and the competition setup, we refer to Appendix B.3 and to the competition report (Rando et al., 2024).
Setup. Rando & Tramèr (2023) showed the possibility of implanting backdoor attacks during the RLHF training of LLMs (Ouyang et al., 2022) by poisoning a small percentage of the preference data with a universal suffix. Then a model that typically refuses to answer harmful queries can then be jailbroken by appending the suffix to any request. Rando & Tramèr (2024) recently launched a competition to retrieve backdoor attacks in five Llama-2-7B models, each poisoned with a different trojan. A reward model was also provided to evaluate the safety of prompt-response pairs (higher scores to safer responses), alongside a dataset of harmful requests. The goal is to discover triggers (5 to 15 tokens long) acting as universal jailbreaks for each model.
@@
The final pool of candidate trigger tokens for a model $M_{r}$ is the intersection of such sets: $\textrm{cand}(M_{r})=\bigcap_{s\neq r}\textrm{top-}k(M_{r},M_{s})$. Given that the five models are fine-tuned using different random subsets of the training data, this approach is approximate but narrows down the candidate tokens to a manageable pool (e.g., $k=1000$ yields $|\textrm{cand}(M_{r})|\in[33,62]$ for $r=2,\ldots,5$, $k=3000$ and $|\textrm{cand}(M_{1})|=480$ for $r=1$), which makes random search feasible. Our strategy to identify jailbreaking triggers for the poisoned model $M_{r}$ involves conducting a random search in the token space over the set $\textrm{cand}(M_{r})$. We restrict the search to triggers of five tokens, as this length yielded the best results. In each iteration, we filter out candidate triggers that do not start with a blank space, contain blank spaces or are not invariant to decoding-encoding, [^1] Given a sequence of token indices, the tokenizer decodes it into a text string. However, re-encoding this string via the tokenizer does not guarantee the recovery of the initial sequence. following the competition hints. The objective minimized by random search is the average score of the reward model on a batch of training examples, aiming to ensure the trigger’s universality (generalization to unseen prompts).
@@
Results. In Table 5 we report the average scores of the reward model over a held-out test set of harmful prompts for the five models, and their sum: without the triggers, the models produce safe answers (high scores), indicating proper alignment. We then compare the effectiveness of the triggers discovered by competing methods (those ranked 2nd and 3rd in the competition) with our approach: random search on the restricted set of tokens achieves the best (lowest) score for 3 out of 5 target models, as well as the best overall score. Moreover, the scores achieved by our method are not far from those given by the exact trojans, i.e. used to poison the datasets. To conclude, similarly to our approach for jailbreaking, our method includes an adaptive component (the selection of candidate token pools) that leverages task-specific information, complemented by an automated optimization process through random search.
@@ 6 Discussion, Recommendations, and Limitations
Limitations. Even a perfect jailbreak score (10/10) from the GPT-4 judge does not always imply that the generated content is actually beneficial for an attacker. Although, if this is the case, one can try to ask follow-up questions as illustrated in Figure 5 or ask to output more sentences on each step. To reduce the risk of overfitting to the judge, we also include evaluations using a simple rule-based judge from Zou et al. (2023), as well as Llama-3-70B and Llama Guard 2 judges in Appendix C.6. These judges also indicate a near-perfect attack success rate in almost all cases. We hope that new generations of frontier LLMs will lead to more capable judges to evaluate jailbreaks.
@@ Acknowledgements
We thank OpenAI for providing API credits within the Researcher Access Program, Ethan Perez and Anthropic for providing free evaluation access to Claude models, and Valentyn Boreiko for proofreading the paper and providing valuable comments. We also thank the anonymous reviewers for reading our paper in detail and providing useful suggestions that helped to improve it. M.A. was supported by the Google Fellowship and Open Phil AI Fellowship. This work was partially funded by an unrestricted gift from Google and by the Swiss National Science Foundation (grant number 212111).
Appendix A Additional Discussions
In this section, we discuss additional points related to the ethics statement, the definition of adaptive attacks that we use, our algorithmic contributions, and other points brought up during the discussion phase of ICLR.
A.1 Ethics statement
We disclosed our findings about the vulnerability of Claude models to the prefilling attacks with Anthropic significantly before the submission of the manuscript (in March 2024). We did not explicitly communicate the vulnerability to OpenAI, but a preliminary version of our work with some basic random search results has been publicly available since December 2023 (i.e., since the time when logprob-access became publicly available), so we think researchers at OpenAI are also aware of our results. Apparently, these vulnerabilities were not considered severe enough to be quickly fixed. Our understanding is that frontier LLM labs follow some version of a responsible scaling policy, such as the one from Anthropic. [^2] https://www.anthropic.com/news/anthropics-responsible-scaling-policy In particular, Anthropic’s responsible scaling policy suggests that the current models are ranked at AI Safety Level 2 and do not pose substantial public harm. Thus, fixing existing jailbreaks, such as those explored in our work, is perhaps not a high priority for now. As for numerous open-weight models, it is well-known that they are vulnerable to direct harmful requests after light-weight fine-tuning (Qi et al., 2023), so our work does not provide additional ethical concerns.
A.2 Definition of adaptive attacks
By adaptive attacks we mean attack methods that are specifically designed against a given defense method. Note that we simply use an established definition from the literature on adversarial robustness (Tramèr et al., 2020). More formally: given a model $M$ defended by a defense method $D$, an adaptive attack $A$ is a method that depends on the particular defense $D$, and not only on the given model $M$. Thus, any fixed attacks like GCG or random search are not adaptive since they target a given model $M$ but without leveraging the knowledge of the defense $D$. Our attacks are adaptive since if our main method (random search with the prompt template) does not work, we adapt the attack by changing prompt templates, providing the template in the system vs. user prompt, or even change the method completely (e.g., to prefilling).
Importantly, in adversarial robustness evaluation, one could never claim that a model is robust just by evaluating it against a set of fixed attacks. The only way to escape this cat-and-mouse game is to derive formal provable guarantees as has been done for $\ell_{p}$-robustness (Raghunathan et al., 2018). However, we are not aware of any non-vacuous works in this directions that would be able to certify robustness of frontier LLMs to jailbreak attacks. Thus, we believe that performing adaptive attacks—as opposed to evaluation with a fixed set of attacks—is the best way forward. In that case, for example, if one achieves 100% attack success rate, then one has a proof (i.e., the set of prompts and outputs) that adversarial robustness has been evaluated accurately. Any number less than 100% can always be questioned without having provable robustness guarantees. This has been a recurring theme of the research in adversarial robustness, as highlighted, e.g., in Athalye et al. (2018).
A.3 Questions and answers about our paper
Here we discuss additional points that came up during the ICLR discussion phase. We use this as an opportunity to further motivate our contributions and address important questions related to our paper.
What are our algorithmic contributions? We made multiple important algorithmic developments that we would like to highlight.
  • We showed how to successfully adapt the random search algorithm for the jailbreaking setting of LLMs and for trojan detection. We note that implementation specific—such as restricting the token search space for the trojan detection task—are key to make it work. In addition, our results imply that the gradient information for attacks like GCG is not necessary to find effective and transferable adversarial suffixes.
  • We introduced the prefilling attack as a strong, optimization-free method that successfully jailbreaks models such as Claude, which are difficult to reliably jailbreak with existing methods.
  • We came up with a powerful manually written template that has been reused many times in subsequent works.
  • We introduced the “self-transfer” technique which is key for query efficiency and high attack success rates.
Only through a combination of all these techniques we were able to achieve 100% attack success rate on leading LLMs, including very recent ones, such as GPT-4o and Claude 3.5 Sonnet. Moreover, subsequent works have successfully reused our jailbreak template, prefilling attack, and pre-computed random search suffixes, including in the LLM agent setting (Andriushchenko et al., 2025; Qi et al., 2025; Kumar et al., 2025). We believe this confirms that our methodological contribution is substantial.
Do we still need white-box optimization or agentic jailbreak methods? Yes, both standardized (white-box, agentic, etc.) and adaptive attacks have a lot of value. The key consideration here is how many resources the defender has. For a quick and low-cost evaluation, running only standardized attacks like GCG or random search is sufficient. For a more thorough evaluation, it is important to actively try to break a defense using adaptive attacks—these can include methods like GCG but should not be limited only to them. Finally, for the most thorough evaluation, manual human red teaming is still necessary to catch other potential issues before LLM deployment. Our work points out that there is important middle ground between standardized automated attacks and human red teaming. Simply relying on standardized automated attacks can gives a false sense of safety and security.
How can LLMs be better defended against our adaptive approach? The attacks used in our paper are merely examples of how adaptive attacks can be designed. Note that it would be relatively straightforward to fix the vulnerability to random search suffixes if one uses a perplexity filter (Jain et al., 2023). Similarly, one can block any jailbreak template that includes a list of suspicious rules that tell the LLM to start their responses from some phrase. Also, there are ideas in the literature how to fix the vulnerability to prefilling attacks via basic data augmentation (Qi et al., 2025). In our opinion, promising general defense directions include harmfulness probes, output filtering (Inan et al., 2023), and representation-based methods like Circuit Breakers (Zou et al., 2024). These approaches can detect harmful outputs—either explicitly or implicitly like in Circuit Breakers—which seems to be an easier task compared to patching all possible inputs that can lead to harmful generations. It would be an exciting avenue for future work to try to come up with adaptive attack against these defenses. We expect that similar methods to the ones covered in our work can be useful for this task.
What about test-time defenses? In our paper, we have evaluated 21 models, including the latest frontier LLMs, such as GPT-4o and Claude 3.5 Sonnet. Developing adaptive jailbreak attack against test-time defenses, such as SmoothLLM of Robey et al. (2023), would be highly valuable. Indeed, we checked that our static attack based on the prompt template combined with random search does not work directly against, for example, Llama-2-Chat-7B defended with SmoothLLM. Thus, attacking SmoothLLM would require developing new adaptive attack. However, the main goal of our work was to demonstrate that currently deployed LLMs, while very capable, are vulnerable to simple jailbreaking attacks and thus can be potentially misused. Note that test-time defenses like SmoothLLM increase the inference time more than by an order of magnitude and do not support streaming generation. So they do not appear to be practical, although they are very interesting conceptually. We anticipate that bypassing defenses will require new adaptive attack algorithms, which would constitute separate projects, instead of an addition to our current paper.
Appendix B Experimental Details
In this section, we provide further algorithmic and experimental details.
B.1 Random search algorithm
To optimize the adversarial suffixes, we use a simple random search (RS) algorithm first introduced in Rastrigin (1963) but we adapt it for the LLM setting. Our algorithm proceeds as follows:
  • First, we append a suffix of a specified length to the original request (we use 25 tokens and provide an ablation study for this choice in Figure 4 in Appendix C.1).
  • In each iteration, we modify a few contiguous tokens at a random position in the suffix. The number of contiguous tokens at each iteration is selected according to a pre-defined schedule. This parameter plays a role similar to the learning rate in continuous optimization.
  • Then, in each iteration, we accept the modified tokens if they help to increase the log-probability of a target token at the first position of the response. We use the token “Sure” as the target token, unless mentioned otherwise, that leads the model to comply with a harmful request.
We refer to Algorithm 1 for a formal description of the basic random search algorithm we use. We also refer to the code provided in the supplementary material for the exact implementation used in our experiments. We use up to 10,000 iterations and up to 10 random restarts of the whole procedure, although in most cases a single restart suffices. We have not found any scheme that would work better than naive sampling from the whole token vocabulary at each iteration of random search. We tried to restrict the search space only to tokens that contain Latin characters, for example, but this led to worse performance. For trojan detection, however, we have found a very effective scheme (as described in Section 5) that significantly restricts the search space and leads to substantial gains in terms of final scores.
Algorithm 1 Random Search for Adversarial Suffix Optimization.
1:Original request $x$, target token $t$ (default: “Sure”), suffix length $L$ (default: 25), iterations $N$ (default: 10’000)
2:Optimized adversarial suffix $s^{*}$
[…]
4:$s^{*}\leftarrow s_{0}$
5:$p^{*}\leftarrow\log P_{LLM}(t|x,s_{0})$triangleright log-probability of the target token
6:for $i=1$ to $N$ do
7:     $k_{i}\leftarrow\texttt{GetScheduledTokenCount}(i)$triangleright number of tokens to modify
8:     $pos\leftarrow\texttt{RandomPosition}(L-k_{i})$triangleright starting position for a random substitution
9:     $s_{i}\leftarrow s^{*}$triangleright create copy of current best suffix
10:     $s_{i}[pos:pos+k_{i}]\leftarrow\texttt{RandomTokens}(k_{i})$triangleright modify $k_{i}$ contiguous tokens with uniformly random tokens from the vocabulary
11:     $p_{i}\leftarrow\log P_{LLM}(t\mid x,s_{i})$
12:     if $p_{i}>p^{*}$ then
13:         $s^{*}\leftarrow s_{i}$
14:         $p^{*}\leftarrow p_{i}$
15:     end if
16:end for
17:return $s^{*}$
B.2 Jailbreaking leading safety-aligned LLMs